OpenStudy (anonymous):

will medal please explain have 4 like this Isabel has a job transporting soft drinks by truck. Her truck is filled with cans that weigh 14 ounces each and bottles that weigh 70 ounces each. There is a combined total of 950 cans and bottles in her truck. Let x be the number of 14-ounce cans in her truck. Write an expression for the combined total weight (in ounces) of the cans and bottles in her truck.

OpenStudy (anonymous):

If we let x be cans, then let y represent bottles

OpenStudy (anonymous):

hence the total weight would be the weight of how many bottles she has + the weight of how many cans she has

OpenStudy (anonymous):

so say x represents bottles, how much does x bottles weigh?

OpenStudy (anonymous):

17x right?

OpenStudy (anonymous):

because 1 bottles weighs 17ounces (i.e when x=1... and so on)

OpenStudy (anonymous):

sorry, x is cans not bottles. sorry for any confusion

OpenStudy (anonymous):

oops, cans weigh14 ounces not 17 haha my typing is poor

OpenStudy (anonymous):

its ok i figured what u meant

OpenStudy (anonymous):

so x=14

OpenStudy (anonymous):

no, x is a can

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

14ounces is how much one can weighs

OpenStudy (anonymous):

so how much would 2 cans weigh (i.e x=2)...it would be 14*2=28ounces right?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so how much would x cans weigh and y bottles weigh

OpenStudy (anonymous):

so x cans would weigh 14x and y bottles would weigh 70y

OpenStudy (anonymous):

hence, total weight = 14x+70y if we combine the bottles and cans right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good so we also know that x+y=950 (no. of cans and bottles) and total weight=14x+70y

OpenStudy (anonymous):

seems here we aren't told how many bottles and cans of each there are, only that there are 950 cans and bottles together

OpenStudy (anonymous):

im confused lol x = 14 ounce cans bottles are 70 ounces 950 total Write an expression for the combined total weight (in ounces) of the cans and bottles in her truck.

OpenStudy (anonymous):

no x does not equal 14

OpenStudy (anonymous):

14 represents HOW MUCH x weighs

OpenStudy (anonymous):

so we write it as 14x

OpenStudy (anonymous):

yes u are sorry that came out wrong

OpenStudy (anonymous):

because if i said i was holding two cans, that means x=2 which also means 2 cans weighs, 14x=14*x=28 ounces

OpenStudy (anonymous):

14x+70(950-x) ???

OpenStudy (anonymous):

that irrelevent to be honest, our expression is simply: Total weight = 14x+70y

OpenStudy (anonymous):

it seems that 950 cans and bottles is redundant for the total weight expression unless we know how many cans and bottles there are

OpenStudy (anonymous):

thanks but i inputted that answer and it was correct

OpenStudy (anonymous):

haha interesting... that only gives an expression for the weight of bottles haha

OpenStudy (anonymous):

im just being pedantic thats all. the problem statement needs re-wording if thats the answer

OpenStudy (anonymous):

but well done!